# How Physics and Background Attributes Impact Video Transformers in Robotic Manipulation: A Case Study on Planar Pushing

### v3

@@ V-C Experiments with Varying Physics Attributes
Color. We compare the zero-shot performance of models trained on data with a single green Ball or red Ball. The GP values for this group are the highest among four controlled experiments, illustrating that VTs are highly sensitive to object color. This can be explained by the fact that shifts in object color dramatically affect pixel value of the input videos thus causing a significant perturbation in the inputs of the VOT models. Pure appearance factors as simple as color which should not affect trajectory prediction in the robotic manipulation context hence can still cause significant adverse effects on VT’s performance.
Friction. As mentioned before, Foam and Ball share near identical shape and color but Ball features in lower friction coefficient compared with Foam. Therefore, we compare the performance of this group to understand the impact of friction. When the force on the gripper is large enough to overcome the static friction, the observed effect of the differing friction coefficients is that objects with lower rolling friction coefficient, such as Ball, starts rolling on the plane and is bounced back from the workspace boundary, resulting in longer and more unpredictable trajectories. We observe asymmetries on the GPs of VOT-MaxViT and VOT-MaxViT-2, which reveal that it may be easier for models trained on longer trajectories to generalize. However, this conclusion doesn’t hold true for VOT-Swin-T. This could potentially result from the different architectures between MaxViT [62] and Swin Transformer [5]. The Swin Transformer in particular has a more local focus with strong inductive bias, complicating the prediction of a fast-moving object (Ball).
Shape. Here we consider the effect of changing object shapes with a comparative study of Cube and Icosahedron, which are both 3D printed from identical material. Based on the average GPs across three controlled physics attribute groups, it appears that color has a greater impact than shape on performance, while shape and friction show a more comparable level of influence on GPs.
@@ V-D Fine-Tuning and Qualitative Observations
Figure 4: PE vs fine-tuning dataset size graphs. Models are pre-trained for 90 epochs on Single dataset and then fine-tuned on Quintuple data for 10 epochs and evaluated on Quintuple. The x-axis represents the size of the fine-tuning datasets. (a) Fine-tuning curves of VOT-MaxViT (VOT-1) and VOT-MaxViT-2 (VOT-2) with x-axis ranging from 300 to 3000. (b) Fine-tuning curves of VOT-Swin-T (SWIN) with x-axis ranging from 700 to 7000.
@@
Figure 5: Examples of predicted and corresponding ground-truth trajectories. (a) A complicated long trajectory of an object colliding with the boundary four times. (b) A successful prediction that featured heavy occlusion in the top camera view. (c) A failure mode. Note that the discontinuities in labels and predicted trajectories are a result from the downsampling of input videos.
In Fig.5, we present examples of predicted and ground truth trajectories - while VTs are able to predict complex long (a) or occluded (b) trajectories, failure cases also present in the case of (c). In (c), the target object (Ball) exhibits an interesting unusual oscillation motion pattern possibly due to scratches on the plexiglass plate, representing a case not seen during training. Videos of examples described above and data saturation curve of VOT-MaxViT can be found in the supplementary video.
@@ V-E Limitations and Conclusions
While we have presented empirical results backed by a large real-world dataset, various factors such as the down-sampling of frame rate and resolution that was necessary for computational complexity reasons may affect our observations. Furthermore, we focus on typical VT architecture design components in the specific narrow setting of planar pushing in order to be able to report results at large data scale. Our experiments across the data subsets provide empirical evidence that: i) When training from scratch, VTs’ performance appears to decline with increasing background scene complexity. Furthermore, models trained on dynamic backgrounds exhibit enhanced generalisability as compared to static backgrounds in zero-shot scenarios. ii) Among the physics attributes considered (color, friction coefficient, shape), the VTs’ performance displays the greatest sensitivity to color - which is not a design goal in the context of robotic manipulation. Notably, shape and color demonstrate symmetry in cross zero-shot generalisability tests, while friction exhibits asymmetry, potentially attributed to model architecture. iii) Regarding the fine-tuning process, a fixed required fine-tuning dataset size across settings isn’t observed. It depends on the specific model structure and the fine-tuning can even have negative effect on model performance if the dataset size is not enough.

### v4

@@ V-C Experiments with Varying Physics Attributes
Color. We compare the zero-shot performance of models trained on data with a single green Ball or red Ball. The GP values for this group are the highest among the four controlled experiments, illustrating that VTs are highly sensitive to object color. This can be explained by the fact that shifts in object color dramatically affect pixel value of the input videos thus causing a significant perturbation in the inputs of the VOT models. Pure appearance factors as simple as color which should not affect trajectory prediction in the robotic manipulation context hence can still cause significant adverse effects on VT’s performance.
Friction. As mentioned before, Foam and Ball share near identical shape and color but Ball features a lower friction coefficient compared with Foam. Therefore, we compare the performance of this group to understand the impact of friction. When the force on the gripper is large enough to overcome the static friction, the observed effect of the differing friction coefficients is that objects with lower rolling friction coefficient, such as Ball, start rolling on the plane and are bounced back from the workspace boundary, resulting in longer and more unpredictable trajectories. We observe asymmetries on the GPs of VOT-MaxViT and VOT-MaxViT-2, which reveal that it may be easier for models trained on longer trajectories to generalize. However, this conclusion does not hold true for VOT-Swin-T. This could potentially result from the different architectures between MaxViT [62] and Swin Transformer [5]. The Swin Transformer in particular has a more local focus with strong inductive bias, complicating the prediction of a fast-moving object (Ball).
Shape. Here we consider the effect of changing object shapes with a comparative study of Cube and Icosahedron, which are both 3D printed from the same plastic material. Based on the average GPs across three controlled physics attribute groups, it appears that color has a greater impact than shape on performance, while shape and friction show a more comparable level of influence on GPs.
@@ V-D Fine-Tuning and Qualitative Observations
Figure 4: PE vs fine-tuning dataset size graphs. Models are pre-trained for 90 epochs on Single dataset and then fine-tuned on Quintuple data for 10 epochs and evaluated on Quintuple. The x-axis represents the number of the fine-tuning videos. (a) Fine-tuning curves of VOT-MaxViT (VOT-1) and VOT-MaxViT-2 (VOT-2) with x-axis ranging from 300 to 3000. (b) Fine-tuning curves of VOT-Swin-T (SWIN) with x-axis ranging from 700 to 7000.
@@
Figure 5: Examples of predicted and corresponding ground-truth trajectories, with Ball being the target object. (a) A complicated long trajectory of an object colliding with the boundary four times. (b) A successful prediction that featured heavy occlusion in the top camera view. (c) A failure mode. Note that the discontinuities in labels and predicted trajectories are a result of the downsampling of input videos. Refer to the supplementary video for more details.
In Fig.5, we present examples of predicted and ground truth trajectories - while VTs are able to predict complex long (a) or occluded (b) trajectories, failure cases also present in the case of (c). In (c), the target object (Ball) exhibits an interesting unusual oscillation motion pattern possibly due to scratches on the plexiglass plate, representing a case not seen during training. Videos of examples described above and a data saturation curve of VOT-MaxViT can be found in the supplementary video.
@@ V-E Limitations and Conclusions
While we have presented empirical results backed by a large real-world dataset, various factors such as the down-sampling of frame rate and resolution that was necessary for computational complexity reasons may affect our observations. Furthermore, we focus on typical VT architecture design components in the specific narrow setting of planar pushing in order to be able to report results at large data scales. Our experiments across the data subsets provide empirical evidence that: i) When training from scratch, VTs’ performance appears to decline with increasing background scene complexity. Furthermore, models trained on dynamic backgrounds exhibit enhanced generalisability as compared to static backgrounds in zero-shot scenarios. ii) Among the physics attributes considered (color, friction coefficient, shape), the VTs’ performance displays the greatest sensitivity to color - which is not a design goal in the context of robotic manipulation. Notably, shape and color demonstrate symmetry in cross zero-shot generalisability tests, while friction exhibits asymmetry, potentially due to the model architecture. iii) Regarding the fine-tuning process, we observed variations in optimal fine-tuning dataset sizes across settings and models. Furthermore, we observe that fine-tuning may even have a negative effect on model performance if the dataset size is not large enough.